OpenStudy (jasbe):

What is the inverse of g(x) = x + 8 / 3 ? G - 1 (x) = 3x + 8 G - 1 ( x) = x + 3 / 8 G - 1 ( x) = 3x - 8 G - 1 (x) = 3x + 3/-8

jimthompson5910 (jim_thompson5910):

The given function is this right? \[\Large g(x) = \frac{x+8}{3}\]

OpenStudy (jasbe):

@jim_thompson5910 yes

jimthompson5910 (jim_thompson5910):

So we follow these steps step 1) Replace g(x) with y step 2) Swap x and y step 3) solve for y \[\Large g(x) = \frac{x+8}{3}\] \[\Large y = \frac{x+8}{3}\] \[\Large x = \frac{y+8}{3}\] are you able to solve for y?

OpenStudy (jasbe):

Am I subtracting 3 since its positive ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the first step is to multiply both sides by 3 this is to undo the division of 3

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{y+8}{3}\] \[\Large 3x = y+8\] what comes next?

OpenStudy (jasbe):

You subtract y ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

nope you subtract 8 from both sides \[\Large 3x = y+8\] \[\Large 3x-8 = y+8-8\] \[\Large 3x-8 = y\] \[\Large y = 3x-8\] So the inverse function is \[\Large g^{-1}(x) = 3x-8\]

OpenStudy (jasbe):

Ok thank you @jim_thompson5910